OpenStudy (anonymous):

how to do division of hexadecimal numbers? like 57F6D by B9?

OpenStudy (anonymous):

@shevron

OpenStudy (anonymous):

it should give you 79b

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

Dividing numbers work the same for any base

OpenStudy (anonymous):

i wanna know the method

OpenStudy (anonymous):

I just showed you the method. Dividing hexadecimal numbers uses exactly the same method as dividing decimal numbers.

OpenStudy (anonymous):

@wiseworld, did you look at the picture (snap.jpg) I attached? It shows the long division steps of the hex example you gave.

OpenStudy (anonymous):

i dint understand

OpenStudy (anonymous):

In decimal, the rightmost digit is the "ones" place. The next digit is the "tens" place The next is "hundreds", and so on. This is the same as the rightmost digit is the 10^1 place, the next is 10^2, the next is 10^3, and so on. In hex, the rightmost digit is also the ones place. The next digit is the "twohundredfiftysixes" place The next digit is the "fourthousandninetysixes" place and so on. This is because the rightmost digit is 16^1, the next is 16^2, the next is 16^3, and so on. This is the basis for carrying and borrowing when you manipulate numbers mathematically, regardless of the number base. Am I explaining this to you or am I confusing you more?

OpenStudy (anonymous):

@wiseworld Do you understand now, or do you want me to try explaining further?

OpenStudy (anonymous):

You could also put in the decimal point and expand this out to 79b.92ae5b54692...

OpenStudy (queelius):

I think it would make a lot more sense for most people to first convert the base-16 numbers to base-10 instead of trying to do all of the long-division arithmetic in base-16. Of course it's not necessary, but that'll be a lot more intuitive for most people.

OpenStudy (queelius):

So: B9 in base 10: 9 + 11*16 = 185 57F6D = 13 + 16*6 + 16^2*15 + 16^3*7 + 16^4 * 5 = 360301 So, 360301 / 185 = 1947 remainder 106.

OpenStudy (anonymous):

Whether it's simpler or more intuitive or not isn't the issue. What he asked was how to do division using hexadecimal numbers.

OpenStudy (queelius):

He said "how to do division of hexadecimal numbers" ... I don't think that necessarily means he requires a long division process that only uses hexademical numbers. So, anyway, I hope my answer was helpful to him.